Please help me.
What kind of a formula can I apply to it?
The diameter of the planet Mercury is approximately 3032 miles. Find the distance around its equator.(Use 3.14 to approximate π.)

Answer by user_dude2008(1862)

diameter: d=3032

C=pi*d=3.14*3032=9520.48


Answer: 9,520.48 miles